### v1

# Multimodal Large Language Models to Support Real-World Fact-Checking

Abstract
Multimodal large language models (MLLMs) carry the potential to support humans in processing vast amounts of information. While MLLMs are already being used as a fact-checking tool, their abilities and limitations in this regard are understudied. Here is aim to bridge this gap. In particular, we propose a framework for systematically assessing the capacity of current multimodal models to facilitate real-world fact-checking. Our methodology is evidence-free, leveraging only these models’ intrinsic knowledge and reasoning capabilities. By designing prompts that extract models’ predictions, explanations, and confidence levels, we delve into research questions concerning model accuracy, robustness, and reasons for failure. We empirically find that (1) GPT-4V exhibits superior performance in identifying malicious and misleading multimodal claims, with the ability to explain the unreasonable aspects and underlying motives, and (2) existing open-source models exhibit strong biases and are highly sensitive to the prompt. Our study offers insights into combating false multimodal information and building secure, trustworthy multimodal models. To the best of our knowledge, we are the first to evaluate MLLMs for real-world fact-checking.
1 Introduction
Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Mubashara et al. (2023). The visual component in these cases could be manipulated, used out-of-context, or accompanied by text that makes a false claim Huh et al. (2018); Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). In such cases, fact-checkers and the tools they employ need to be able to handle multiple modalities.
Large language models (LLMs) store extensive information about the world, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al. (2023); Hu et al. (2023a). As such, they can be a powerful tool in the hands of fact-checkers, who often seek extra factual knowledge to verify a claim Nakov et al. (2021). The application of LLMs to fact-checking has been extensively studied in a text-only setting, wherein a model receives as input both a textual claim, with or without a set of evidence documents, and needs to reason to determine the veracity of the claim Chen et al. (2022); Pan et al. (2023); Li et al. (2023a). LLMs have increasingly strong understanding and reasoning capabilities, so using them to sift through evidence documents when verifying a claim is proving very effective Chen et al. (2022); Li et al. (2023a). Yet, this pipeline approach requires an extra step of evidence retrieval, which is computationally heavy and error-prone, while it overlooks the vast amount of knowledge stored in the parameters of LLMs Hu et al. (2023b). Additionally, LLMs are now giving way from text-only to multimodal models, which can process signals from multiple sources and perform a wider range of tasks Liu et al. (2023b); Zhang et al. (2023); Yin et al. (2023). We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for retrieving evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
Despite their great promise, however, MLLMs have limitations as well, specifically with respect to factuality Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023). Models may produce hallucinations Huang et al. (2023); Li et al. (2023b) and suffer from a lack of knowledge or exhibit biases Hu et al. (2023a); Wang et al. (2023). To understand MLLMs’ potential to support fact-checking real-world claims, we propose the evaluation framework illustrated in Figure 1. It involves querying models for predictions, explanations, and confidence scores on multimodal check-worthy claims. Based on the data obtained, we can further assess the models’ capabilities from various perspectives, addressing the following research questions:
RQ1: How accurate are MLLMs in identifying misinformation in multimodal claims?
RQ2: How do different modalities (text and image) contribute to the assessment of multimodal content?
RQ3: Are models sensitive to prompts in different languages or expressions, or do they exhibit robustness?
RQ4: Can in-context learning enhance model performance?
RQ5: Can MLLMs accurately estimate the correctness of their responses in terms of a confidence score?
RQ6: What types of errors are most common for MLLMs employed as multimodal fact-checkers?
Our investigation spans an array of vision-language models, including GPT-4V, LLaVA-1.5 Liu et al. (2023b), MiniGPT-v2 Chen et al. (2023), and InstructBLIP Dai et al. (2023), which stand as leaders in this domain. We evaluate these models across three existing multimodal fact-checking datasets Zlatkova et al. (2019); Aneja et al. (2023); Yao et al. (2023), and one new dataset, which we build from recent check-worthy claims to mitigate the risk of data contamination.
Our study finds that MLLMs display a range of fact-checking capabilities. GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration, with the accuracy on different datasets generally reaching around . On the other hand, open-source models like MiniGPT-v2 and InstructBLIP lack explanatory and uncertainty-reporting capabilities. LLaVA models show a strong sensitivity to prompt formulation, a tendency to comply with the question in the prompt, and hallucinatory behaviors.
In summary, we systematically evaluate existing multimodal models to determine their capability to support real-world fact-checking. We formulate several basic research questions and assess their accuracy, robustness, calibration, and reasons for failure across various datasets. Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on both fact-checking and MLLMs’ factuality. As far as we are aware, our study is the first assessment of MLLMs in the context of real-world fact-checking.
2 Background and Related Work
LLMs for text-only fact-checking
Hoes et al. (2023) find that ChatGPT accurately classifies 69% of statements in a dataset built from the fact-checking platform PolitiFact. FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions. Studies have also focused on leveraging the capabilities of LLMs to develop automated frameworks for decomposing claims, generating queries, and predicting based on gathered evidence, thereby augmenting the interpretability of the fact-checking process Pan et al. (2023); Li et al. (2023a). However, these works often focused solely on prediction accuracy. Our work analyzes the model’s fact-checking capabilities from three perspectives: predictions, explanations, and confidence.
Multimodal fact-checking
Multimodal misinformation has been found to spread faster than its text-only counterparts Li and Xie (2020). It primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). The multimodal fact-checking is gnerally structured into distinct stages within pipelines, including claim detection extraction, evidence retrieval, verdict prediction, and explanation Nakov et al. (2021); Mubashara et al. (2023). In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the targeted claims.
Evaluation of MLLMs
With the advancement of MLLMs, the systematic evaluation of their capabilities has become increasingly critical. A broad array of benchmarks has been established to estimate the MLLM’s performance across a variety of tasks, including perceptual tasks such as image recognition, image description, and optical character recognition, as well as cognitive tasks like commonsense reasoning and numerical calculation Yin et al. (2023); Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023); Tong et al. (2024). Tu et al. (2023) access the resilience of MLLMs against various image-related attacks from a security perspective. However, there is an absence of research on evaluating the ability to conduct multimodal fact-checking with MLLMs. Our work aims to bridge this gap.
3 Evaluation Framework
We propose an evaluation framework that analyzes model capabilities along the various dimensions illustrated in Figure 1. Our experiments are conducted in an evidence-free setting, meaning that models have access only to the multimodal claims without any external knowledge. Through extensive controlled experiments, we conduct a detailed analysis of models’ potential in fact-checking real-world claims, including their accuracy and reliability.
3.1 Evaluation Prompt
We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below. This prompt is formulated following the example of prompts found in related fact-checking work Pan et al. (2023); Min et al. (2023), while for acquiring explanations and confidence estimates we refer to Xiong et al. (2023).
Is it true that "CLAIM"? True or False? Use the following format to provide your answer:
Prediction: [True or False]
Explanation: [put your evidence and step-by-step reasoning here]
Confidence Level: [please show the percentage]
Note: The confidence level indicates the degree of certainty you have about your answer and is represented as a percentage. For instance, if your confidence level is , it means you are certain that your answer is correct and there is a chance that it may be incorrect.
3.2 Metrics
Response Types
Based on the three components (prediction, explanation, and confidence score) in the model’s response, we categorize them into four types: True, False, Uncertain, and Deny. Deny refers to instances where the model fails to respond according to our requirements, missing any component. Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc. Others are classified as True or False based on the model’s stated prediction. This classification scheme is designed to reveal the model’s inherent biases through the distribution of True and False responses, ability to follow instructions the number of Deny responses, and assess its knowledge-awareness through the frequency of Uncertain cases.
Accuracy Metrics
Assuming the total count of True (T), False (F), Uncertain (U) and Deny (D) responses is . Since not all responses include predictions, we denote the number of samples containing predictions as , with . We introduce two distinct accuracy metrics, True/False Accuracy is defined as , and Overall Accuracy is defined as , where represent the number of correctly predicted samples within the corresponding group. The first metric posits that predictions made with expressed uncertainty should be omitted from accuracy assessments. The second metric provides a holistic prediction-based score over the same set of data points for all models, thus allowing for a fair comparison.
3.3 Models
We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview). We also experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al. 2023b, a), InstructBLIP (Flan-T5-xxl, Dai et al. 2023), and MiniGPT (v2, Chen et al. 2023). All parameters are set to the default values, with max_tokens=300 for GPT-4V. These models are notably representative and competitive in a multimodal setup.111We hereafter omit the version numbers of the open-source models, using LLaVA(7b), LLaVA(13b), MiniGPT, and InstructBLIP instead.
3.4 Datasets
Fauxtography |
Cosmos |
MOCHEG |
Post-4V |
True |
200 |
240 |
267 |
81 |
False |
200 |
240 |
237 |
105 |
Total |
400 |
480 |
504 |
186 |
Fauxtography
is a multimodal fact-checking dataset sourced from the websites Snopes and Reuters Zlatkova et al. (2019). Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
COSMOS
is a dataset built from Snopes and News Outlets, and tailored for training models to autonomously identify image and text pairs that are out of context Aneja et al. (2023), i.e. falsely paired. We randomly select 240 true and 240 false samples from the test split.
MOCHEG
is originally based on textual claims from Snopes and PolitiFact, with associated images serving as evidence Yao et al. (2023). This implies that a single claim can correspond to multiple images, many of which are not explicitly required for the claim to be verified. We first identify multimodal claims through keywords such as photograph, image, etc. Subsequently, we manually select images that were explicitly requested for verification and filter out those with overlaying words like fake, misleading, or miscaptioned. This procedure results in a total of 504 data entries.
Post-4V
To mitigate potential data leakage, we obtain new data from the Snopes based on articles published after the release of GPT-4V (Sept. 26, 2023). We adhere to the same filtering approach used for MOCHEG, resulting in 186 samples.
Details of all data samples used are presented in Table 1. Overall, the sample distribution is balanced, containing only two types of labels: True and False.222We exclude data labeled as Mixture, Unproven, and other labels that lack clear veracity. Additionally, we collected the URLs of the web pages corresponding to the claims and their fact-checking articles.
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Image |
Model |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| Yes |
GPT |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
LLa.(7b) |
337 |
6 |
1 |
56 |
53.4 |
46.3 |
449 |
7 |
4 |
20 |
52.0 |
50.0 |
409 |
6 |
0 |
89 |
54.2 |
44.8 |
157 |
1 |
3 |
25 |
42.0 |
37.1 |
LLa.(13b) |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
MiniGPT |
55.5 |
62.1 |
54.3 |
55.9 |
Instruct. |
54.3 |
74.4 |
52.6 |
59.1 |
| No |
GPT |
105 |
202 |
32 |
61 |
80.1 |
71.0 |
111 |
253 |
50 |
66 |
64.0 |
57.5 |
153 |
265 |
38 |
48 |
73.9 |
68.9 |
35 |
68 |
10 |
73 |
74.8 |
67.7 |
LLa.(7b) |
398 |
0 |
2 |
0 |
50.3 |
50.0 |
474 |
0 |
4 |
2 |
50.2 |
49.8 |
502 |
0 |
2 |
0 |
53.0 |
52.8 |
181 |
0 |
4 |
1 |
43.1 |
43.0 |
LLa.(13b) |
167 |
214 |
17 |
2 |
62.5 |
62.0 |
263 |
195 |
20 |
2 |
76.6 |
75.8 |
168 |
316 |
20 |
0 |
62.0 |
61.5 |
40 |
124 |
22 |
0 |
60.4 |
60.8 |
MiniGPT |
56.8 |
60.4 |
54.0 |
58.1 |
Instruct. |
55.8 |
68.3 |
48.6 |
61.8 |
Yes |
Baseline |
50.0 |
50.0 |
53.0 |
56.5 |
4 Evaluation and Analysis
4.1 RQ1: Comprehensive Performance
Table 2 presents the comprehensive performance of various models across four datasets. The "Image" column clarifies the input type, with "Yes" indicating multimodal inputs and "No" signifying text-only inputs. MiniGPT and InstructBLIP, while incapable of providing explanations and confidence scores, still manage to generate predictions; thus, only the Overall Accuracy metric is assessed. The accuracy of the majority class is noted at the bottom of the table.
The upper section of Table 2 reveals that GPT-4V outperforms other models in both True/False and Overall Accuracy metrics, with True/False Accuracy frequently exceeding Overall Accuracy. It implies GPT-4V’s heightened precision when it responds with confidence. LLaVA’s accuracy consistently improves with increased model size. MiniGPT and InstructBLIP surpass the baseline and LLaVA(7b), nearing the accuracy levels of the 13b model. In addition, LLaVA(7b) shows a strong bias, almost always predicting True. As the model size increases, predictions become more balanced, and the expression of uncertainty rises. Furthermore, there is a significant reduction in the number of Deny responses, indicating that larger models are more capable of following instructions. The high Deny responses from GPT-4V are mainly due to its safety alignment, often replying with statements like "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy."
Figure 2 displays multimodal claims and corresponding responses from GPT-4V and LLaVA(13b), with the annotations w. Image and w/o Image next to the model names indicating whether an image is in the input, respectively. From the perspective of explanations, GPT-4V’s responses contain a richer set of facts. For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case. The 5th example provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates other details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 4th image. In comparison, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 2nd and 4th examples. Additionally, LLaVA’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity, indicating inefficient reasoning.
4.2 RQ2: Impact of Images: Multimodal VS. Text-Only inputs
By comparing data in Table 2, it is evident that images significantly impact the predictions of GPT-4V, with the model’s accuracy decreasing to various extents. Simultaneously, the instances where the model refuses to respond have noticeably increased. As in Figure 2, GPT-4V’s explanations indicate its awareness of missing images, using phrases like "there is no verified public record" or "without any provided context." This suggests that GPT-4V engages in knowledge retrieval to verify the existence or context of the images.
Counterintuitively, open-source models demonstrate higher accuracy with text-only inputs in many cases compared to multimodal inputs. To understand this, we scrutinize their explanations and find two specific scenarios. First, when the claim is false and the image is altered, as illustrated in the 1st and 4th examples in Figure 2, models can accurately predict the veracity of textual claims. However, when images are included, the models, failing to recognize the alterations, are misled into making incorrect predictions. Second, when both the image and the claim are true, as demonstrated in the 3rd example, models that consider images might deem the claim False due to inadequate evidence. In contrast, predictions for textual claims remain true, even with hallucinations highlighted in yellow. This implies that the models were not exposed to enough negative examples during the fine-tuning phase.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
EN |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
ZH |
144 |
188 |
48 |
20 |
78.6 |
75.3 |
180 |
168 |
100 |
32 |
83.0 |
72.5 |
216 |
198 |
50 |
40 |
84.3 |
79.2 |
62 |
82 |
30 |
12 |
78.5 |
68.3 |
DE |
154 |
173 |
30 |
43 |
80.7 |
72.3 |
156 |
223 |
45 |
56 |
77.0 |
67.1 |
224 |
198 |
17 |
65 |
83.9 |
75.0 |
55 |
87 |
17 |
27 |
76.8 |
64.5 |
AR |
101 |
82 |
14 |
203 |
73.8 |
73.5 |
88 |
109 |
19 |
264 |
68.0 |
64.6 |
137 |
77 |
13 |
277 |
84.1 |
78.2 |
29 |
26 |
9 |
122 |
79.6 |
69.4 |
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
idx |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| GPT-4V |
0 |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
2 |
158 |
181 |
32 |
29 |
84.1 |
81.3 |
163 |
219 |
75 |
23 |
85.1 |
80.8 |
205 |
234 |
34 |
31 |
82.1 |
81.2 |
55 |
95 |
22 |
14 |
77.8 |
75.3 |
5 |
143 |
202 |
31 |
24 |
81.2 |
76.3 |
175 |
216 |
72 |
17 |
88.0 |
81.5 |
179 |
250 |
44 |
31 |
78.6 |
75.6 |
51 |
106 |
18 |
11 |
76.4 |
71.5 |
| LLaVA |
0 |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
1 |
197 |
163 |
7 |
33 |
48.6 |
48.5 |
169 |
236 |
15 |
60 |
69.6 |
66.3 |
247 |
211 |
9 |
37 |
57.6 |
57.9 |
68 |
105 |
5 |
8 |
56.6 |
57.5 |
2 |
170 |
223 |
4 |
3 |
49.1 |
49.5 |
197 |
274 |
7 |
2 |
69.0 |
68.5 |
318 |
178 |
7 |
1 |
58.1 |
57.9 |
41 |
141 |
4 |
0 |
56.0 |
57.0 |
3 |
243 |
145 |
8 |
4 |
52.6 |
53.0 |
243 |
222 |
13 |
2 |
71.8 |
70.2 |
295 |
197 |
11 |
1 |
57.3 |
56.3 |
82 |
93 |
10 |
1 |
56.5 |
56.5 |
4 |
211 |
180 |
6 |
3 |
62.1 |
62.0 |
232 |
244 |
3 |
1 |
77.3 |
77.1 |
210 |
286 |
4 |
4 |
57.3 |
56.3 |
55 |
129 |
2 |
0 |
57.6 |
57.0 |
5 |
241 |
154 |
3 |
2 |
63.3 |
63.0 |
313 |
163 |
4 |
0 |
80.5 |
80.4 |
260 |
236 |
8 |
0 |
62.1 |
62.3 |
109 |
75 |
2 |
0 |
52.7 |
52.7 |
4.3 RQ3: Robustness to Prompts
Multilingual prompts
Since our datasets primarily originate in English-language websites, multilingual prompts can assess the adaptability of intrinsic knowledge across different languages and mitigate data leakage. We translate our prompt template and claims into Chinese, German, and Arabic using the Google Translate API. Table 3 presents the performance across different languages. Notable variations can be observed in adherence to instructions, particularly in Arabic, which exhibits a higher frequency of Deny responses across datasets. Nevertheless, the overall accuracy levels remain comparable, suggesting that the responses’ lack of explanations and confidence scores in Arabic did not significantly impact model performance. Moreover, the True/False Accuracy generally surpasses the Overall Accuracy, indicating GPT-4V’s good knowledge awareness. We do not conduct experiments on open-source models as they are predominantly trained and tested in English.
Prompts with similar semantics
We employ ChatGPT to generate semantically similar yet distinct prompts, as depicted in Figure 3. The part pertaining to explanations and confidence scores is identical to the original prompt. We randomly select two prompts for testing GPT-4V and all generated prompts for the LLaVA(13b) model. The outcomes of the experiments are detailed in Table 4, where Index 0 denotes the original prompt, and the other indices represent the newly crafted prompts. The data reveals that GPT-4V maintains commendable consistency, while LLaVA demonstrates variable biases across different prompts. For the same dataset, such as Fauxtography, models tend to predict the majority of claims as True when using the original prompt (index 0). In contrast, the prompt of index 2 leads the models to predict most claims as False. Moreover, on Fauxtography and COSMOS, the maximum discrepancy in overall accuracy reached and , respectively, compared to just and for GPT-4V. It indicates that the LLaVA(13b) model is highly sensitive to prompts and fails to maintain consistency.
4.4 RQ4: Improving Capabilities via In-Context Learning
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Demo |
Num |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
0 |
0 |
286 |
98 |
16 |
0 |
54.4 |
52.8 |
331 |
134 |
14 |
1 |
69.7 |
67.7 |
363 |
112 |
29 |
0 |
61.5 |
58.1 |
116 |
65 |
5 |
0 |
56.4 |
55.9 |
| 1 |
1 |
244 |
135 |
9 |
12 |
60.9 |
61.8 |
295 |
171 |
7 |
7 |
75.3 |
74.2 |
255 |
229 |
12 |
8 |
62.6 |
62.7 |
95 |
88 |
3 |
0 |
56.3 |
55.4 |
2 |
182 |
161 |
6 |
51 |
59.5 |
60.0 |
234 |
220 |
8 |
18 |
75.8 |
73.5 |
169 |
282 |
10 |
43 |
57.4 |
55.5 |
42 |
109 |
6 |
29 |
67.5 |
63.4 |
| 2 |
1 |
264 |
121 |
4 |
11 |
61.6 |
62.8 |
337 |
131 |
7 |
5 |
73.9 |
73.5 |
314 |
180 |
4 |
6 |
64.8 |
65.1 |
97 |
79 |
10 |
0 |
57.4 |
55.9 |
2 |
176 |
204 |
10 |
10 |
60.0 |
60.5 |
233 |
230 |
11 |
6 |
80.3 |
79.6 |
199 |
291 |
8 |
6 |
61.2 |
60.7 |
44 |
131 |
11 |
0 |
60.0 |
60.8 |
| 3 |
1 |
211 |
185 |
1 |
3 |
62.6 |
62.8 |
279 |
195 |
6 |
0 |
72.2 |
71.7 |
236 |
266 |
2 |
0 |
61.8 |
61.5 |
75 |
111 |
0 |
0 |
60.2 |
60.2 |
2 |
191 |
198 |
9 |
2 |
64.1 |
63.8 |
263 |
201 |
16 |
0 |
75.4 |
75.4 |
206 |
288 |
9 |
1 |
61.1 |
61.7 |
54 |
127 |
4 |
1 |
59.7 |
59.7 |
| 4 |
1 |
191 |
193 |
9 |
7 |
62.2 |
62.0 |
260 |
203 |
16 |
1 |
78.0 |
77.9 |
210 |
283 |
11 |
0 |
60.6 |
60.9 |
50 |
126 |
8 |
2 |
60.8 |
60.8 |
2 |
196 |
188 |
7 |
9 |
62.8 |
62.8 |
259 |
207 |
12 |
2 |
77.5 |
77.5 |
207 |
284 |
11 |
2 |
61.1 |
61.3 |
52 |
122 |
10 |
2 |
61.5 |
61.8 |
As demonstrated in the 1st and 4th examples in Figure 2, the LLaVA(13b) model struggles to distinguish claims associated with manipulated images. The image content is consistent with the textual claim but contradicts real-world facts. To enhance the model’s comprehension of the task, checking facts rather than the consistency between the image and text, we experiment with employing in-context learning to improve its performance. In detail, we obtain GPT-4V’s responses for claims outside our test datasets and employ these as instructive examples for LLaVA(13b). Due to LLaVA’s limitation with multiple image inputs, we combine different images vertically with a 20-pixel gap between them. The linkage between images and claims is narratively indicated, as illustrated in the following prompt:
Use the following format to answer whether the claim regarding the pictures is True or False:
SAME FORMAT AS IN THE ORIGINAL PROMPT.
For the first picture, claim: "CLAIM1"
Prediction: False
Explanation: XXX
Confidence Level: XXX
For the second picture, claim: "CLAIM2"
Prediction: True
Explanation: XXX
Confidence Level: XXX
For the third picture, claim: "CLAIM"
We construct four sets of demonstrations, each comprising one true and one false claim. We report the performances when using only the false claim and employing both claims as examples. We observe a universal improvement in model accuracy when examples are presented in Table 5. However, the use of the second example does not contribute to accuracy. In the explanations provided by LLaVA, we can observe two effects: one is the average length of the explanation. For instance, on MOCHEG, with the presentation of 0, 1, and 2 examples from the first set, the average sentence length of the model’s explanations are 3.09, 4.02, and 3.75, respectively. Another change is the frequency of the term "manipulated," which increases from 4 to 170 and 193 on MOCHEG. This suggests that the model can learn to analyze from the perspective of image manipulation. The last two rows of Figure 2 display responses from the LLaVA(13b) model. In the 4th example, it is evident that through learning, the model recognizes the altered part in the image and explains the motive behind the modifications: "but the gun and menacing pose are added elements," and "image is likely created to spread misinformation and discredit Joe Biden or to provoke a reaction from viewers." However, some samples are not correctly identified, such as the 1st example, and we observe that LLaVA produces outdated information, as highlighted in green.
4.5 RQ5: Calibration Evaluation
Calibration refers to the alignment between a model’s predicted probabilities (or confidence levels) and the actual outcomes. It reflects the model’s knowledge-awareness and ability to accurately estimate the confidence in its responses Geng et al. (2023). In this evaluation, we use verbalized confidence, which is readily accessible. Figure 4 displays the confidence distributions alongside calibration curves for GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals. For instance, a value of 95 corresponds to confidence levels within the bracket. GPT-4V’s confidence scores are largely clustered in the interval, whereas LLaVA(13b)’s scores are more frequently found in the range. The calibration curves reveal that both models exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores. Conversely, LLaVA(13b)’s curve suggests a propensity for overconfidence. The Expected Calibration Errors (ECEs, Guo et al. 2017), detailed in Table 6, further substantiate that GPT-4V maintains superior calibration.
ECE |
Fauxtography |
COSMOS |
MOCHEG |
Post-4V |
GPT-4V |
7.76 |
4.36 |
4.88 |
11.83 |
LLaVA |
30.20 |
15.26 |
22.82 |
28.45 |
4.6 RQ6: Taxonomy of Failure Reasons
Upon examining GPT-4V’s explanations, we identify six primary categories of failure reasons. Failures attributed to Lack of Knowledge occur when responses are based on trivial or incomplete information, often contradicting common sense. Recognition Errors arise from the model’s visual module misinterpreting images, including failing to identify altered or AI-generated visuals. Reasoning Errors occur when the model incorrectly interprets images due to factors like unnatural distortions or unusual perspectives, or there are self-contradictions in reasoning. Failures due to Partial Facts happen when the model addresses only a subset of the facts present in a textual claim. Image Issues involve failures related to poor image quality, such as blurriness or incompleteness, affecting the model’s ability to recognize content accurately. Ambiguity Claims involve semantic uncertainties, where a term could reference multiple entities. Examples beyond the lack of knowledge are presented in Figure 5. In subfigure (a), GPT-4V misinterprets scars from a dog’s burn as being painted or digitally altered. In subfigure (b), it correctly identifies the locations of Notre Dame and the Eiffel Tower but engages in erroneous reasoning. In subfigure (c), the model understands that a kiss can release hormones like oxytocin and knows that MR images cannot capture chemical reactions, with its predictions primarily influenced by the first fact. In subfigure (d), the model cannot discern that the scene depicted is related to posture training based on the image’s content. In subfigure (e), "Playboy" has multiple meanings, which could refer to a publisher or magazine’s name. We also manually tall the frequency of these errors across various datasets, as detailed in Table 7. Overall, the primary cause of errors is attributed to Lack of Knowledge, suggesting that augmenting the system with additional evidence could enhance its performance. There is room for improving the model’s ability to understand and reason about AI-generated or distorted images.
Know |
Recog |
Reas |
Partial |
Image |
Ambi |
| Fauxtography |
59.3 |
20.3 |
17.2 |
1.6 |
1.6 |
0.0 |
| COSMOS |
80.3 |
5.9 |
9.8 |
2.0 |
0 |
2.0 |
| MOCHEG |
73.2 |
3.6 |
7.1 |
3.6 |
5.4 |
7.1 |
| Post-4V |
85.2 |
11.1 |
0.0 |
0.0 |
0.0 |
3.7 |
We also explore the utilization of GPT-4 for detecting nonfactual content within explanations when fact-checking articles are available. We design a prompt comprising two steps: removing sentences containing no facts and predicting whether the reference article supports the sentence. Due to the limited content space of the article, we do not present the specific prompt. Figure 6 showcases two instances where, although the predictions are correct, the facts within the explanations are inaccurate. This underscores the necessity for further examination of their veracity. The photograph on the left, featuring a young girl named Fouzia, was captured in Morocco by Giacomo Pirozzi, a photographer for the United Nations Children’s Fund (UNICEF). However, GPT-4V incorrectly attributes it to photographer Terri Gold in Yemen, 1998. The right photo is incorrectly identified as from the movie "Platoon," actually, it is from the movie "Tropic Thunder." The returned sentences also contained substantial image descriptions not present in the original fact-checking articles. Improving the automatic assessment of the factuality in explanations is a focus for our future work.
5 Conclusion and Future Work
We investigated the capabilities of MLLMs to fact-check real-world claims, relying solely on their parametric knowledge and reasoning capabilities without external references. We proposed an evaluation framework, designed various experiments to address the research questions. Our results indicated that state-of-the-art MLLMs, such as GPT-4V, have the potential to assist professional fact-checkers: they can enhance their efficiency by providing reference predictions, valuable clues, and explanatory insights, together with confidence. Open-source models exhibit sensitivity to text, with different prompts revealing various biases. Moreover, they are susceptible to being deceived by altered images, and even in the absence of images, they tend to produce hallucinations pertaining to image descriptions. Our work demonstrates both the potential and limitations of MLLMs, laying the groundwork for further leveraging MLLMs to support fact-checking.
In future work, we plan to investigate how to enhance the fact-checking capabilities of models, focusing not only on improving accuracy but also on strengthening robustness and augmenting the knowledge reasoning and the verification abilities required for fact-checking. We further plan to explore how MLLMs can better support fact-checking when connected to external knowledge.